Coronado Fabricators produces protective covers for smart phones. Since the covers must be customized to each smart phone model, Coronado uses a job order costing system. On September 1, the company reported the following inventory balances:

Direct Materials $26,700
Work in Process $148,000
Finished Goods $265,000

During September, the following events occurred:

1. Coronado purchased direct materials costing $420,000 on account.
2. Coronado used $429,000 in direct materials in production.
3. Coronado’s employees clocked 20,000 direct labor hours at an average wage rate of $11.01 per direct labor hour.
4. The company incurred $173,000 in manufacturing overhead, including $19,000 in indirect labor costs.
5. Using direct labor hours as the application base, the company applied $170,000 of manufacturing overhead to jobs worked on in September.
6. The company completed production on jobs costing $853,000.
7. The company delivered jobs costing $1,079,000 to customers.

1. Calculate the ending September balance of the Direct Materials, Work in Process, and Finished Goods Inventory accounts.

2. Calculate total manufacturing costs for September.
Answer #1

1) Ending balance

Direct material = Beginning+Purchase-Used = 26700+420000-429000 = 17700

Work in process = Beginning+Total manufacturing cost-Cost of goods completed = 148000+429000+220200+170000-853000 = 114200

Finished goods = Beginning+Cost of goods completed-Cost of goods sold = 265000+853000-1079000 = 39000

2) Total manufacturing cost

Direct material 429000
Direct labor 220200
Overhead 170000
Total manufacturing cost 819200